# ROTI-GCV: Generalized Cross-Validation for right-ROTationally Invariant Data

### v2

@@ 3.1.1 Estimators for $r^{2},\sigma^{2}$
Consistent estimation of $\sigma^{2}$ has been studied in [23]. We propose a different approach since we observe it performs better in finite samples. In particular, we establish the following result:
@@
$\frac{1}{n}\sum_{i=1}^{n}\left(y_{i}-\mathbf{x}_{i}^{\top}\hat{\boldsymbol{\beta}}_{\lambda}\right)^{2}-\left[r^{2}\frac{\lambda^{2}}{\gamma}\left(v_{\mathbf{D}}(-\lambda)-\lambda v^{\prime}_{\mathbf{D}}(-\lambda)\right)+\sigma^{2}\lambda^{2}v^{\prime}_{\mathbf{D}}(-\lambda)\right]\xrightarrow{a.s.}0.$ (8)
Observe that Eq. (8) can be used as an estimating equation, meaning that, if one treats (8) as an equality to 0 (i.e. replacing “$\xrightarrow{a.s.}$” with “$=$”), then we can solve for the unknown quantities $r^{2}$ and $\sigma^{2}$.
Recall that $\gamma=p/n$, $v_{\mathbf{D}}$, and $v^{\prime}_{\mathbf{D}}$, defined in Section 2.1, are all data dependent quantities.
Now, Eq. (8) yields a distinct estimating equation for each value of $\lambda$. While it takes only two equations to solve for $r^{2}$ and $\sigma^{2}$, we instead use a grid of $\lambda$ and find OLS estimates for $r^{2}$ and $\sigma^{2}$ to improve robustness. Explicitly, the scheme proceeds as shown in steps 1-6 of Algorithm 1.
@@
Under Assumptions A1-A7 and a non-degeneracy condition on $\mathbf{D}$ (see Appendix A.7.1 for details), $\hat{r}^{2}$ and $\hat{\sigma}^{2}$ from Algorithm 1 are strongly consistent for $r^{2}$ and $\sigma^{2}$, respectively.
@@
1:$\mathbf{X},\mathbf{y}$, $\lambda$; for use in estimation, grid of regularization strengths $\lambda_{1}<\lambda_{2}\cdots<\lambda_{L}$;
@@
8:$\hat{\sigma}^{2}_{\{\lambda_{i}\}_{i=1}^{L}}(\mathbf{X},\mathbf{y})\leftarrow\frac{\sum_{i=1}^{L}\left(\frac{t_{i}}{a_{i}}-\frac{t_{1}}{a_{1}}\right)\left(\frac{b_{i}}{a_{i}}-\frac{b_{1}}{a_{1}}\right)}{\sum_{i=1}^{L}\left(\frac{b_{i}}{a_{i}}-\frac{b_{1}}{a_{1}}\right)^{2}}$,   $\hat{r}^{2}_{\{\lambda_{i}\}_{i=1}^{L}}(\mathbf{X},\mathbf{y})\leftarrow\frac{\sum_{i=1}^{L}\left(\frac{t_{i}}{b_{i}}-\frac{t_{1}}{b_{1}}\right)\left(\frac{a_{i}}{b_{i}}-\frac{a_{1}}{b_{1}}\right)}{\sum_{i=1}^{L}\left(\frac{a_{i}}{b_{i}}-\frac{a_{1}}{b_{1}}\right)^{2}}.$
@@ 3.1.2 Plug-in estimation
Uniform convergence, unlike pointwise convergence, ensures the risk attained by the minimizer (over a compact interval) of our cross-validation metric is close to optimal. See Appendix B.4 for a short proof.
@@ 4 GCV under Signal-PC Alignment
An issue that emerges when attempting to apply ROTI-GCV in practice is that the signal often tends to align with the top eigenvectors of the data, violating the independence assumption imposed on $\boldsymbol{\beta}$ and $\mathbf{O}$. The right-rotationally invariant assumption, for fixed $\boldsymbol{\beta}$, inherently assumes that the signal is incoherent with respect to the eigenbasis. However, this assumption is generally not true, and in fact in the i.i.d. anisotropic case, the geometry of $\boldsymbol{\beta}$ with respect to the top eigenvectors of $\boldsymbol{\Sigma}$ is known to significantly influence the behavior of ridge regression (see e.g. [42]).
Adapting the approach given in [23], we model this scenario using two index sets $\mathcal{J}_{a}$ and $\mathcal{J}_{c}$ of distinguished eigenvectors, which are aligned and coupled eigenvectors, respectively. We replace Assumptions A6 and A7 in Section 3 with the following:
@@
    The true signal $\boldsymbol{\beta}$ is given as $\boldsymbol{\beta}=\boldsymbol{\beta}^{\prime}+\sum_{i\in\mathcal{J}_{a}}\sqrt{n}\alpha_{i}\mathbf{o}_{i}$, where $\|\boldsymbol{\beta}^{\prime}\|/\sqrt{n}=r$; $\boldsymbol{\beta}^{\prime}$ is independent of $(\mathbf{X},\boldsymbol{\epsilon})$. An eigenvector $\mathbf{o}_{i}$ is an aligned eigenvector if $\mathbf{o}_{i}\in\mathcal{J}_{a}$, as then $\mathbf{o}_{i}$ aligns with the signal $\boldsymbol{\beta}$;
@@ 2. A11
    lim $\limsup\|\mathbb{E}[\widetilde{\mathbf{D}}^{\top}\widetilde{\mathbf{D}}]\|_{\mathsf{op}}<C$ for some constant $C$. This differs from Assumption A2, as it is on the test data.
The main difficulty that arises in our setting, compared to [23], is that we must account for how the geometry of the eigenvectors of the test set relates to those of the training set. One should expect, for example, in some types of structured data that the top eigenvectors of the training data and test data are closely aligned. We account for this effect through the coupled eigenvectors condition (Assumption A9). Furthermore, note that without this coupling, any alignment between the eigenvectors of the training set and the signal $\boldsymbol{\beta}$ will not exist in the test set, since then the eigenvectors of the test set would be independent of this alignment. We are still interested in the same test risk $R_{\mathbf{X},\mathbf{y}}$ as in Eq. (2), except now note that $\widetilde{\mathbf{X}}$ and $\mathbf{X}$ are not independent.
@@
$\mathcal{B}_{\mathbf{X}}(\hat{\boldsymbol{\beta}}_{\lambda},\boldsymbol{\beta})$ $=\frac{\lambda^{2}}{n}\sum_{i=1}^{p}\left[\frac{\left(r^{2}/\gamma+\mathds{1}(i\in\mathcal{J}_{a})n\alpha_{i}^{2}\right)}{(D_{ii}^{2}+\lambda)^{2}}\left(\mathfrak{d}_{\mathsf{bulk}}^{2}+\mathds{1}(i\in\mathcal{J}_{c})(\mathfrak{d}_{i}^{2}-\mathfrak{d}^{2}_{\mathsf{bulk}})\right)\right]$
@@ 4.1 Estimating $r^{2}$, $\sigma^{2}$, and $\alpha_{i}$
Our approach is simply to estimate the $\alpha_{i}$’s using the classical principal components regression (PCR);
see [20, 19] for details, also described in steps 1 and 2 of Algorithm 2. We then compute a transformed model which is right-rotationally invariant, from which Algorithm 1 steps 1-6 can be used to find $r^{2}$ and $\sigma^{2}$.
@@
Under the additional assumption that $\boldsymbol{\epsilon}$ has Gaussian entries, $\tilde{r}^{2}$, $\tilde{\sigma}^{2}$, and $\hat{\boldsymbol{\alpha}}$, from Algorithm 2, are all strongly consistent.
The assumption of Gaussian entries is technical, and we believe it can be removed with some work.
@@ 4.2 Cross-validation metric
Our approach for cross-validation is to then again explicitly compute the risk formulas $\mathcal{B}_{\mathbf{X}}(\hat{\boldsymbol{\beta}}_{\lambda},\boldsymbol{\beta})$ and $\mathcal{V}_{\mathbf{X}}(\hat{\boldsymbol{\beta}}_{\lambda},\boldsymbol{\beta})$, where we estimate the unknown parameters from data. To apply our method, we therefore need to consistently estimate $r^{2},\sigma^{2},\boldsymbol{\alpha}$, and $\{\mathfrak{d}_{i}^{2}\}_{i\in\mathcal{J}_{c}}$ and $\mathfrak{d}_{\mathsf{bulk}}^{2}$. Lemma 4 gives us most of these parameters, but we still require consistent estimates of the $\mathfrak{d}_{i}$’s. This is a nonissue in cases where $\mathbf{D}$ and $\widetilde{\mathbf{D}}$ are modeled as deterministic. If they are random, we expect that if the test data is drawn from the same right-rotationally invariant ensemble as the training data, then, in practice, one can use the eigenvalues of the training data to do this. However, our framework further allows for the test distribution to be drawn from a different right-rotationally invariant ensemble. In cases where test data covariates are available (we do not require access to the labels), it is possible to then use test data to do this estimation; this is done for experiments in Section 5. We therefore make one final assumption:
@@ 1. A11
    We assume there exists estimators $\hat{\mathfrak{d}}_{i}$ and $\hat{\mathfrak{d}}_{\mathsf{bulk}}$ where we have $\sup_{i\in\mathcal{J}_{a}\cup\mathcal{J}_{c}}|\hat{\mathfrak{d}}_{i}-\mathfrak{d}_{i}|\xrightarrow{a.s.}0$, and $|\hat{\mathfrak{d}}_{\mathsf{bulk}}-\mathfrak{d}_{\mathsf{bulk}}|\xrightarrow{a.s.}0$.
We then take our validation metric as $\mathsf{aROTI}\text{-}\mathsf{GCV}(\lambda)=\mathcal{R}_{\mathbf{X},\mathbf{y}}(\tilde{r}^{2},\tilde{\sigma}^{2},\hat{\boldsymbol{\alpha}},\{\hat{\mathfrak{d}}_{i}\}_{i\in\mathcal{J}_{c},\mathsf{bulk}})$, where the latter expression denotes substituting every parameter with its estimated version. Algorithm 2 summarizes our entire procedure.
@@
1:$\mathbf{X},\mathbf{y}$, regularization strength $\lambda$; index sets $\mathcal{J}_{a},\mathcal{J}_{c}$; consistent estimators $\hat{\mathfrak{d}}_{i}^{2}$ and $\hat{\mathfrak{d}}_{\mathsf{bulk}}^{2}$; for use in estimation of $r^{2},\sigma^{2}$, grid of regularization strengths $\lambda_{1},\dots,\lambda_{L}$;
@@
3:$\mathbf{X}_{\mathsf{proj}}\leftarrow\mathbf{X}\mathbf{O}_{\mathcal{J}_{a}}^{\top}$
4:$\hat{\boldsymbol{\alpha}}\leftarrow(\mathbf{X}_{\mathsf{proj}}^{\top}\mathbf{X}_{\mathsf{proj}})^{-1}\mathbf{X}_{\mathsf{proj}}^{\top}\mathbf{y}/\sqrt{n}$
5:$\mathbf{y}_{\mathsf{new}}=(\mathbf{D}_{\bar{\mathcal{J}_{a}}}^{\top}\mathbf{D}_{\bar{\mathcal{J}_{a}}})^{1/2}(\mathbf{X}_{\mathsf{resid}}^{\top}\mathbf{X}_{\mathsf{resid}})^{-1}\mathbf{X}_{\mathsf{resid}}^{\top}\mathbf{y}$
6:$\mathbf{X}_{\mathsf{new}}=(\mathbf{D}_{\bar{\mathcal{J}_{a}}}^{\top}\mathbf{D}_{\bar{\mathcal{J}_{a}}})^{1/2}\mathbf{O}_{\bar{\mathcal{J}_{a}}}$
7:Compute $\tilde{r}=\hat{r}^{2}_{\{\lambda_{i}\}_{i=1}^{L}}(\mathbf{X}_{\mathsf{new}},\mathbf{y}_{\mathsf{new}}),\tilde{\sigma}^{2}=\hat{\sigma}^{2}_{\{\lambda_{i}\}_{i=1}^{L}}(\mathbf{X}_{\mathsf{new}},\mathbf{y}_{\mathsf{new}})$
@@
This states that aROTI-GCV is uniformly consistent on compact intervals, again ensuring that the minimizer of aROTI-GCV attains risk that is close to optimal.
On selecting $\mathcal{J}_{a},\mathcal{J}_{c}$:
[23] presents a way to do hypothesis testing for alignment to detect $\mathcal{J}_{a}$. For detecting $\mathcal{J}_{c}$, one can measure the angle among the top eigenvectors of the test set and top eigenvectors of the training set, placing pairs of vectors which have overlap far from $90^{\circ}$ into $\mathcal{J}_{c}$. We leave the development of a rigorous hypothesis test to future works.
@@ 5.1 Experiments on right-rotationally invariant data
Figure 1: Risk curves produced by the cross-validation methods in addition to the true risk curve. Each plot uses a different form of right-rotationally invariant data, as indicated in each caption. (b): Autocorrelated data: rows of $\mathbf{X}$ are drawn according to $\mathbf{x}_{i}=\rho\mathbf{x}_{i-1}+\sqrt{1-\rho^{2}}\mathbf{z}_{i}$, with $\mathbf{z}_{i}$ being i.i.d. draws from $\mathsf{N}(0,\mathbf{I}_{n})$. We set $\rho=0.8$. (c): $t$-distributed data: each row is drawn from a multivariate $t$ distribution with $3$ degrees of freedom. (d): Equicorrelated data $\mathbf{X}\in\mathbb{R}^{n\times p}$ has independent columns, but each column follows a multivariate Gaussian distribution with covariance matrix $\boldsymbol{\Sigma}$, where $\Sigma_{ij}=\rho$ if $i\neq j$, and $\Sigma_{ii}=1$. All simulations have $n=p=1000$ and $r^{2}=\sigma^{2}=1$. The $x$-axis for every plot is $\lambda$, the regularization parameter. The colored lines (blue, green, red, purple) are one of 10 iterations. In each iteration, we compute the cross-validation metric over a range of $\lambda$ to produce the line, which reflects the estimated out-of-sample risk. The black line of each plot shows the average result for that method. The dashed blue line shows the average expected MSE curve as a benchmark.
Figure 1 illustrates the performance of ROTI-GCV and compares it to LOOCV and GCV. In these figures, the Tuned Risk (abbreviated TR) refers to the out-of-sample risk obtained when we use the value of $\lambda$ that optimizes the given cross-validation method. Minimum risk (MR), denotes the minimum of the true expected out-of-sample risk curve. Standard errors are shown in parentheses for all tuned risks. Estimated Risk (ER) refers to the value of the out-of-sample risk that would be estimated using the given cross-validation method. As expected, in the Gaussian setting all three methods perform well. However, for dependent data, such as in the autocorrelated and equicorrelated cases, we observe that the risk curves produced by GCV and LOOCV are wildly off from the truth, whereas ROTI-GCV accurately captures the true risk curve. Surprisingly, the tuned risk values from LOOCV, GCV are still decently close to that of ROTI-GCV.
As discussed in Remark 3, GCV (and also LOOCV) estimate a quantity with an additional factor of $\sigma^{2}$; thus, to make the evaluation of estimated risks fair, these two curves are plotted with this term removed.

### v3

@@ 3.1.1 Estimators for $r^{2},\sigma^{2}$
Consistent estimation of $\sigma^{2}$ has been studied in Li and Sur, (2023). We propose a different approach since we observe it performs better in finite samples. In particular, we establish the following result:
@@
[…]
Observe that Eq. (8) can be used as an estimating equation, meaning that, if one treats (8) as an equality to 0 (i.e. replacing $\xrightarrow{a.s.}$” with “==”), then we can solve for the unknown quantities $r^{2}$ and $\sigma^{2}$. Recall that $\gamma=p/n$, $v_{\mathbf{D}}$, and $v^{\prime}_{\mathbf{D}}$, defined in Section 2.1, are all data dependent quantities. Now, Eq. (8) yields a distinct estimating equation for each value of $\lambda$. While it takes only two equations to solve for $r^{2}$ and $\sigma^{2}$, we instead use a grid of $\lambda$ and find OLS estimates for $r^{2}$ and $\sigma^{2}$ to improve robustness. Explicitly, the scheme proceeds as shown in steps 1-6 of Algorithm 1.
@@
Under Assumptions A1-A7 and a non-degeneracy condition on $\mathbf{D}$ (see Appendix 7.7.1 for details), $\hat{r}^{2}$ and $\hat{\sigma}^{2}$ from Algorithm 1 are strongly consistent for $r^{2}$ and $\sigma^{2}$, respectively.
@@
[…]
@@
[…]
@@ 3.1.2 Plug-in estimation
Uniform convergence, unlike pointwise convergence, ensures the risk attained by the minimizer (over a compact interval) of our cross-validation metric is close to optimal. See Appendix 8.4 for a short proof.
@@ 4 GCV UNDER SIGNAL-PC ALIGNMENT
An issue that emerges when attempting to apply ROTI-GCV in practice is that the signal often tends to align with the top eigenvectors of the data, violating the independence assumption imposed on $\boldsymbol{\beta}$ and $\mathbf{O}$. The right-rotationally invariant assumption, for fixed $\boldsymbol{\beta}$, inherently assumes that the signal is incoherent with respect to the eigenbasis. However, this assumption is generally not true, and in fact in the i.i.d. anisotropic case, the geometry of $\boldsymbol{\beta}$ with respect to the top eigenvectors of $\boldsymbol{\Sigma}$ is known to significantly influence the behavior of ridge regression (see e.g. Wu and Xu, (2020)).
Adapting the approach given in Li and Sur, (2023), we model this scenario using two index sets $\mathcal{J}_{a}$ and $\mathcal{J}_{c}$ of distinguished eigenvectors, which are aligned and coupled eigenvectors, respectively. We replace Assumptions A6 and A7 in Section 3 with the following:
@@
    The true signal $\boldsymbol{\beta}$ is given as $\boldsymbol{\beta}=\boldsymbol{\beta}^{\prime}+\sum_{i\in\mathcal{J}_{a}}\sqrt{n}\alpha_{i}\mathbf{o}_{i}$, where $\|\boldsymbol{\beta}^{\prime}\|/\sqrt{n}=r$; $\boldsymbol{\beta}^{\prime}$ is independent of $(\mathbf{X},\boldsymbol{\epsilon})$. Here $\mathbf{o}_{i}$ refers to the $i$th row of $\mathbf{O}$, which is an eigenvector of $\mathbf{X}^{\top}\mathbf{X}$. An eigenvector $\mathbf{o}_{i}$ is an aligned eigenvector if $\mathbf{o}_{i}\in\mathcal{J}_{a}$, as then $\mathbf{o}_{i}$ aligns with the signal $\boldsymbol{\beta}$;
@@ 2. A11
    $\limsup\|\mathbb{E}[\widetilde{\mathbf{D}}^{\top}\widetilde{\mathbf{D}}]\|_{\mathsf{op}}<C$ for some constant $C$. This differs from Assumption A2, as it is on the test data.
The main difficulty that arises in our setting, compared to Li and Sur, (2023), is that we must account for how the geometry of the eigenvectors of the test set relates to those of the training set. One should expect, for example, in some types of structured data that the top eigenvectors of the training data and test data are closely aligned. We account for this through the coupled eigenvectors condition (A9). Furthermore, without this coupling, any alignment between the eigenvectors of the training set and the signal $\boldsymbol{\beta}$ will not exist in the test set, since then the eigenvectors of the test set would be independent of this alignment. We are still interested in the same test risk $R_{\mathbf{X},\mathbf{y}}$ as in Eq. (2), except now note that $\widetilde{\mathbf{X}}$ and $\mathbf{X}$ are dependent.
@@
$\mathcal{B}_{\mathbf{X}}(\hat{\boldsymbol{\beta}}_{\lambda},\boldsymbol{\beta})$ $=\frac{\lambda^{2}}{n}\sum_{i=1}^{p}\left[\frac{\left(r^{2}/\gamma+\mathds{1}(i\in\mathcal{J}_{a})n\alpha_{i}^{2}\right)}{(D_{ii}^{2}+\lambda)^{2}}\right$.
[…]
@@ 4.1 Estimating $r^{2}$, $\sigma^{2}$, and $\alpha_{i}$
Our approach is simply to estimate the $\alpha_{i}$’s using the classical principal components regression (PCR); see Jolliffe, (1982); Hubert and Verboven, (2003) for details, also described in steps 1 and 2 of Algorithm 2. We then compute a transformed model which is right-rotationally invariant, from which Algorithm 1 steps 1-6 can be used to find $r^{2}$ and $\sigma^{2}$.
@@
Under the additional assumption that $\boldsymbol{\epsilon}$ has Gaussian entries [^2] The assumption of Gaussian entries is technical, and we believe it can be removed with some work., $\tilde{r}^{2}$, $\tilde{\sigma}^{2}$, and $\hat{\boldsymbol{\alpha}}$, from Algorithm 2, are all strongly consistent.
@@ 4.2 Cross-validation metric
Our approach for cross-validation is to then again explicitly compute the risk formulas $\mathcal{B}_{\mathbf{X}}(\hat{\boldsymbol{\beta}}_{\lambda},\boldsymbol{\beta})$ and $\mathcal{V}_{\mathbf{X}}(\hat{\boldsymbol{\beta}}_{\lambda},\boldsymbol{\beta})$, where we estimate the unknown parameters from data. To apply our method, we therefore need to consistently estimate $r^{2},\sigma^{2},\boldsymbol{\alpha}$, and $\{\mathfrak{d}_{i}^{2}\}_{i\in\mathcal{J}_{c}}$ and $\mathfrak{d}_{\mathsf{bulk}}^{2}$. Lemma 4 gives us most of these parameters, but we still require consistent estimates of the $\mathfrak{d}_{i}$’s. This is a nonissue in cases where $\mathbf{D}$ and $\widetilde{\mathbf{D}}$ are modeled as deterministic. If they are random, we expect that if the test data is drawn from the same right-rotationally invariant ensemble as the training data, then, in practice, one can use the eigenvalues of the training data to do this. However, our framework further allows for the test distribution to be drawn from a different right-rotationally invariant ensemble. In such cases, it is possible to then use test data to do this estimation because it only depends on the test covariates $\widetilde{\mathbf{X}}$ (which the practitioner may have access to), but not its labels; this is done for experiments in Section 5. We therefore make one final assumption:
@@ 1. A11
    We assume there exist estimators $\hat{\mathfrak{d}}_{i}$ and $\hat{\mathfrak{d}}_{\mathsf{bulk}}$ where we have $\sup_{i\in\mathcal{J}_{a}\cup\mathcal{J}_{c}}|\hat{\mathfrak{d}}_{i}-\mathfrak{d}_{i}|\xrightarrow{a.s.}0$, and $|\hat{\mathfrak{d}}_{\mathsf{bulk}}-\mathfrak{d}_{\mathsf{bulk}}|\xrightarrow{a.s.}0$.
We then take our validation metric as $\mathsf{aROTI}\text{-}\mathsf{GCV}(\lambda)=\mathcal{R}_{\mathbf{X},\mathbf{y}}(\tilde{r}^{2},\tilde{\sigma}^{2},\hat{\boldsymbol{\alpha}},\{\hat{\mathfrak{d}}_{i}\}_{i\in\mathcal{J}_{c},\mathsf{bulk}})$, where the latter expression denotes substituting every parameter with its estimated version. Algorithm 2 summarizes our entire procedure [^3] The algorithm requires a choice of $\mathcal{J}_{a},\mathcal{J}_{c}$; we give a brief overview of selecting $\mathcal{J}_{a}$ and $\mathcal{J}_{c}$ in Section 5.3, followed by a detailed discussion in Appendix 9..
@@
[…]
@@
[…]
@@
This states that aROTI-GCV is uniformly consistent on compact intervals, ensuring the minimizer of aROTI-GCV attains risk that is close to optimal.
@@ 5.1 Experiments on right-rotationally invariant data
Figure 1: Risk curves produced by the cross-validation methods in addition to the true risk curve. Each plot uses a different form of right-rotationally invariant data, as indicated in each caption. (1(b)): Autocorrelated data: rows of $\mathbf{X}$ are drawn according to $\mathbf{x}_{i}=\rho\mathbf{x}_{i-1}+\sqrt{1-\rho^{2}}\mathbf{z}_{i}$, with $\mathbf{z}_{i}$ being i.i.d. draws from $\mathsf{N}(0,\mathbf{I}_{n})$. We set $\rho=0.8$. (1(c)): $t$-distributed data: each row is drawn from a multivariate $t$ distribution with $3$ degrees of freedom. (1(d)): Equicorrelated data $\mathbf{X}\in\mathbb{R}^{n\times p}$ has independent columns, but each column follows a multivariate Gaussian distribution with covariance matrix $\boldsymbol{\Sigma}$, where $\Sigma_{ij}=\rho$ if $i\neq j$, and $\Sigma_{ii}=1$. All simulations have $n=p=1000$ and $r^{2}=\sigma^{2}=1$. The $x$-axis for every plot is $\lambda$, the regularization parameter. The colored lines (blue, green, red, purple) are one of 10 iterations. In each iteration, we compute the cross-validation metric over a range of $\lambda$ to produce the line, which reflects the estimated out-of-sample risk. The black line of each plot shows the average result for that method. The dashed blue line shows the average expected MSE curve as a benchmark.
Figure 1 illustrates the performance of ROTI-GCV and compares it to LOOCV and GCV. In these figures, the Tuned Risk (abbreviated TR) refers to the out-of-sample risk obtained when we use the value of $\lambda$ that optimizes the given cross-validation method. Minimum risk (MR), denotes the minimum of the true expected out-of-sample risk curve. Standard errors are shown in parentheses for all tuned risks. Estimated Risk (ER) refers to the value of the out-of-sample risk that would be estimated using the given cross-validation method. As expected, in the Gaussian setting all three methods perform well. However, for dependent data, such as in the autocorrelated and equicorrelated cases, we observe that the risk curves produced by GCV and LOOCV are wildly off from the truth, whereas ROTI-GCV accurately captures the true risk curve. Surprisingly, the tuned risk values from LOOCV, GCV are still decently close to that of ROTI-GCV. As discussed in Remark 3, GCV (and also LOOCV) estimate a quantity with an additional factor of $\sigma^{2}$; thus, to make the evaluation of estimated risks fair, these two curves are plotted with this term removed.